OpenStudy (anonymous):

Find the x and y intercept. -4x+5y=-20

OpenStudy (angela210793):

1-st u take x=0 and find y,5y=-20--->y=-4 A(0;-4) 2nd-u take y=0 and find x,-4x=-20--->x=5 B(5;0)

OpenStudy (anonymous):

you can find your x intercept by letting y=0 and you can find your y intercept by letting x=0 -4x+5y=-20 here let y =0 we get -4x+5(0)=-20 -4x =-20 x= -20/-4 x= 5 answer -4x+5y=-20 here let x =0 we get -4(0)+5y=-20 5y=-20 y=-20/5=4 answer

OpenStudy (anonymous):

angela is right

OpenStudy (anonymous):

ok thats what i got for the first part its the second part that i am suck on

OpenStudy (angela210793):

y=-4 mark

OpenStudy (angela210793):

what don't u understand Bell?

OpenStudy (anonymous):

hope you understand my solutions hhhahaha ooppss sorry ... y=-4 answer

OpenStudy (anonymous):

wats the second part now

OpenStudy (anonymous):

my example shows that y= -4

OpenStudy (anonymous):

heres how i solved it

OpenStudy (angela210793):

@Vinit u didn't explain my q..:( Am i supposed to try wht a must be 4 all real no.s? :P

OpenStudy (angela210793):

Good job :)

OpenStudy (anonymous):

-4x+5y=-20 4x=20 20\4=5 x=5

OpenStudy (anonymous):

ok guys i think you can handle them now...good job for every one,...thnkssss...

OpenStudy (anonymous):

its y that i am trying to solve for

OpenStudy (anonymous):

wat question dear

OpenStudy (angela210793):

the 1 I posted yday in this case Bella u have y=0 cause this is how u've found x

OpenStudy (anonymous):

-5y=20 do i divide here of multiply

OpenStudy (angela210793):

where did u get -5y=0??

OpenStudy (angela210793):

-5y=20*

OpenStudy (anonymous):

i divide both sides by 5

OpenStudy (angela210793):

u have 5y=-20 and u u divide both parts by 5--->5y/5=-20/5-->y=-4

OpenStudy (anonymous):

oh ok am so stupid i do to the other what i did to the first